Lori Cook has developed the following forecasting​ model: ModifyingAbove y with caret ​= 36.0 ​+ 4.50​x, where ModifyingAbove y with caret ​= demand for Kool Air conditioners and x​ = the outside temperature​ (degrees ​Fahrenheit) ​a) When the temperature outside is 70degrees ​F, demand forecast​ = ??? air conditioners ​(enter your response as an​ integer).
Answer #1

Solution:

y = 36.0 + 4.50x

where, y = demand for Kool Air conditioners

x = the outside temperature (degrees Fahrenheit)

(a) When the temperature outside is 70 degrees F, demand forecast is calculated by putting the value of x = 70 in the above equation.

y = 36.0 + 4.50x

y = 36.0 + (4.50 x 70)

y = 351

Demand forecast = 351 air conditioners